OpenStudy (anonymous):

Suppose you want to invest $500 into a bank account that earns 3.4% compounding weekly. After 4 years how much money will you have in your account?

OpenStudy (anonymous):

$573 was the answer I got

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) plug in P = 500 r = 0.034 n = 52 (there are 52 weeks in a year) t = 4 to get A = 500(1+0.034/52)^(52*4) I'll let you evaluate that

OpenStudy (anonymous):

572.82 when rounded thank you so much!! @jim_thompson5910 could you help me with one question similar or even a formula would help!

OpenStudy (anonymous):

suppose you own a car that is worth 10,000. It is depreciating by a rate of 15%. How long will it take for your car to be worth 4,000? Round you answer to the nearest tenth. I know it is about 6 years, but I am not sure how to find the answer properly in order to find it to the nearest tenth

jimthompson5910 (jim_thompson5910):

use the formula FV = PV*(1-r)^t where, FV = final value PV = present value r = rate of depreciation (how much as a percentage it decreases every year) t = time (in years) In this case, FV = 4000 PV = 10000 r = 0.15 t = unknown (solving for it) so this means FV = PV*(1-r)^t 4000 = 10000*(1-0.15)^t now solve for t

OpenStudy (anonymous):

am I going to have to use a log?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

any time you need to isolate an exponent, you use logs

OpenStudy (anonymous):

ohh okay I think that is where I was getting stuck

OpenStudy (anonymous):

I am getting 9.2.. but that is way off from my original thought!

OpenStudy (anonymous):

so sorry I am new with logarithms and they are throwing me off.

jimthompson5910 (jim_thompson5910):

4000 = 10000*(1-0.15)^t 4000/10000 = (1-0.15)^t 0.4 = (1-0.15)^t log(0.4) = log((1-0.15)^t) log(0.4) = t*log(1-0.15) log(0.4) = t*log(0.85) log(0.4)/log(0.85) = t t = log(0.4)/log(0.85) t = 5.63805542348653 <--- Use a calculator here t = 6 ... round to the nearest whole number

OpenStudy (anonymous):

wow thank you. I did not realize that you write 4000 as .4 etc. I really appreciate it.

jimthompson5910 (jim_thompson5910):

no you write 4000/10000 as 0.4

jimthompson5910 (jim_thompson5910):

not 4000 as 0.4

OpenStudy (anonymous):

well.. I feel very uneducated at the moment haha. thanks for putting it into simple terms for me because apparently I am way off my game today..

jimthompson5910 (jim_thompson5910):

that's ok, just keep practicing and you'll do better

OpenStudy (anonymous):

thank you I am trying to work through a few more examples to see if I got the hang of now that you put it that way

jimthompson5910 (jim_thompson5910):

yw, that's a good idea